OpenStudy (anonymous):

A conducting loops of wire has an area of 250 cm^2 and a resistance of 40 ohm. There is a magnetic field of 200T perpendicular to the loop. At what rate must this field be reduced to induce a current o f.2A in the loop?

OpenStudy (nali):

your could use the equation \[\large V= - N \Delta (\phi B)/\Delta (t) = -N (B A \cos \theta)/ t = - N A \cos (\theta) (\Delta B/\Delta(t)) \]

OpenStudy (nali):

First find the voltage using Voltage = Current times Resistance or V= I times R Then plug in the V and find the ratio of B/t (magnetic field over time)

OpenStudy (nali):

@Grazes

OpenStudy (anonymous):

I have v=8V. Now what?

OpenStudy (nali):

do you mean 80 V ?

OpenStudy (anonymous):

V=IR=(.2A)(40ohm)=8V

OpenStudy (nali):

oh sorry did not notice the point in front of the 2 there

OpenStudy (nali):

anyway, then you plug in V and the area after changing it to m and the magnetic field

OpenStudy (anonymous):

8V=−N(2.5m^2)(ΔB/Δ(t)) Now what?

OpenStudy (nali):



OpenStudy (nali):

cos 90 = 1 so what you have is right so far

OpenStudy (anonymous):

cos90=0.... perpendicular means cos0

OpenStudy (anonymous):

But I got it, thanks.

OpenStudy (nali):

oh yeah sorry again but anyway that is good you got what I did

OpenStudy (nali):

by the way, the question is asking for the ratio, so after finding the time you still need to divide magnetic field by the time to get the ratio

OpenStudy (anonymous):

the ratio is just B/t, isn't it?

OpenStudy (anonymous):

It's one term, no need to divide.

OpenStudy (nali):

no because you do not have the time (exact amount of seconds) so you need to find it first and then you could find the ratio